# MGE-LDM: Joint Latent Diffusion for Simultaneous Music Generation and Source Extraction

### v2

@@ Abstract
[…]
@@ 5.1 Music Generation
The $\mathcal{T}_{3}$ model, trained on MUSDB18 ($\textbf{M}_{u}$) and MoisesDB ($\textbf{M}_{o}$), both containing real recordings rather than synthesized audio, achieves strong performance on their respective test sets. Despite the domain shift, it also performs competitively on Slakh2100, with results comparable to MSDM, highlighting the model’s robustness to cross-domain generalization. Finally, our model trained on the combined dataset comprising lakh2100, MUSDB18, and MoisesDB (denoted as $\mathcal{T}_{4}$), achieves the best overall results on both $\textbf{M}_{u}$ and $\textbf{M}_{o}$. It also outperforms $\mathcal{T}_{3}$ on $\textbf{S}_{\text{Full}}$, benefiting from the broader training distribution.
@@ 6 Limitations
While MGE-LDM provides a flexible, class-agnostic framework for multi-track audio modeling, several limitations remain. First, all experiments are conducted using 16 kHz monaural audio, which constrains upper-frequency resolution and omits spatial cues, thereby limiting realism for high-fidelity or stereo music applications. Second, the model relies on CLAP-based semantic conditioning, which introduces a modality gap between text and audio [99]. This can occasionally lead to semantic drift during extraction
@@
Finally, our model is trained using triplets $(\textit{mix},\textit{submix},\textit{source})$ that satisfy $\textit{mix}=\textit{submix}+\textit{source}$ in waveform space; however, the latent diffusion process does not enforce an explicit additivity constraint for generated triplets. We believe this omission contributes directly to hallucination penomena, where the model extracts source absent from the mixture. Postolache et al. [27] addresses a related issue by enforcing additivity in a discrete VQ-VAQE latent space, estimating the joint likelihood of two sources by counting codebook co-occurrences — effectively modeling $p(z_{\text{mix}}|z_{\text{src}_{1}},z_{\text{src}_{2}})$, where $z_{*}$ are quantized latent codes. Our current pipeline, however, operates in a continuous latent space, which precludes the direct use of such discrete bin-counting methods. Adapting this latent-domain likelihood formulation to continuous spaces, for example, by designing suitable regularizers or adopting a VQ-VAE-based encoder with discrete diffusion [100, 101, 102] or MaskGiT [65]-style generation, represents a promising direction for future work.
@@ Appendix A Background
In this section, we review two foundational components that support our method: (1) the v-objective score-matching formulation for diffusion models [87], which underpins our latent modeling, and (2) the canonical inpainting algorithm introduced by Song et al. [11] and (3) further refined by Lugmayr et al. [89].
@@ A.3 RePaint
Lugmayr et al. proposed RePaint [89], a resampling-based mechanism that improves score-based inpainting by repeating the diffusion process across multiple forward–reverse cycles. Their key insight is that, in conventional inpainting (as described in Equation (21)), the sampled noise for the known region is independent of the generated (inpainted) region. This lack of synchronization can lead to semantic inconsistencies and disharmony between the known and unknown parts of the sample.
@@ Appendix B Adaptive Timestep Conditioning and Score Approximation
where the approximation assumes that $\textbf{z}_{0}^{\Gamma}\sim q(\mathbf{x}_{0}^{\Gamma}|\textbf{z}_{0}^{\Omega})$ is available from the dataset. Unlike the marginalization-based approximation in Eq (21), this expression introduces no sampling noise during inference, thereby reducing variance.
@@
$\nabla_{\textbf{z}_{\tau}^{\Gamma}}\log q_{\tau}(\textbf{z}_{\tau}^{\Gamma}|\textbf{z}_{0}^{\Omega})$ $\approx\frac{\alpha_{\tau}\textbf{z}_{0}^{\Gamma}-\textbf{z}_{\tau}^{\Gamma}}{\beta^{2}}$ (25)
$\approx\frac{\alpha_{\tau}\hat{\textbf{z}}_{\theta}(\textbf{z}_{\tau}^{\Gamma},\textbf{z}_{0}^{\Omega},\boldsymbol{\tau}^{\Gamma})-\textbf{z}_{\tau}^{\Gamma}}{\beta^{2}},$ (26)

### v3

@@ Abstract
We present MGE-LDM, a unified latent diffusion framework for simultaneous music generation, source imputation, and query-driven source separation. Unlike prior approaches constrained to fixed instrument classes, MGE-LDM learns a joint distribution over full mixtures, submixtures, and individual stems within a single compact latent diffusion model. At inference, MGE-LDM enables (1) complete mixture generation, (2) partial generation (i.e., source imputation), and (3) text-conditioned extraction of arbitrary sources. By formulating both separation and imputation as conditional inpainting tasks in the latent space, our approach supports flexible, class-agnostic manipulation of arbitrary instrument sources. Notably, MGE-LDM can be trained jointly across heterogeneous multi-track datasets (e.g., Slakh2100, MUSDB18, MoisesDB) without relying on predefined instrument categories. Audio samples are available at our project page $\dagger$.
@@ 5.1 Music Generation
The $\mathcal{T}_{3}$ model, trained on MUSDB18 ($\textbf{M}_{u}$) and MoisesDB ($\textbf{M}_{o}$), both containing real recordings rather than synthesized audio, achieves strong performance on their respective test sets. Despite the domain shift, it also performs competitively on Slakh2100, with results comparable to MSDM, highlighting the model’s robustness to cross-domain generalization. Finally, our model trained on the combined dataset comprising Slakh2100, MUSDB18, and MoisesDB (denoted as $\mathcal{T}_{4}$), achieves the best overall results on both $\textbf{M}_{u}$ and $\textbf{M}_{o}$. It also outperforms $\mathcal{T}_{3}$ on $\textbf{S}_{\text{Full}}$, benefiting from the broader training distribution.
@@ 6 Limitations
While MGE-LDM provides a flexible, class-agnostic framework for multi-track audio modeling, several limitations remain. First, all experiments are conducted using 16 kHz monaural audio, which constrains upper-frequency resolution and omits spatial cues, thereby limiting realism for high-fidelity or stereo music applications. Second, the model relies on CLAP-based semantic conditioning, which introduces a modality gap between text and audio [99]. This can occasionally lead to semantic drift during extraction, resulting in irrelevant or hallucinated sources, particularly for stems with limited training data.
@@
Finally, our model is trained using triplets $(\textit{mix},\textit{submix},\textit{source})$ that satisfy $\textit{mix}=\textit{submix}+\textit{source}$ in waveform space; however, the latent diffusion process does not enforce an explicit additivity constraint for generated triplets. We believe this omission contributes directly to hallucination phenomena, where the model extracts sources absent from the mixture. Postolache et al. [27] addresses a related issue by enforcing additivity in a discrete VQ-VAE latent space, estimating the joint likelihood of two sources by counting codebook co-occurrences — effectively modeling $p(z_{\text{mix}}|z_{\text{src}_{1}},z_{\text{src}_{2}})$, where $z_{*}$ are quantized latent codes. Our current pipeline, however, operates in a continuous latent space, which precludes the direct use of such discrete bin-counting methods. Adapting this latent-domain likelihood formulation to continuous spaces, for example, by designing suitable regularizers or adopting a VQ-VAE-based encoder with discrete diffusion [100, 101, 102] or MaskGiT [65]-style generation, represents a promising direction for future work.
@@ Appendix A Background
In this section, we review three foundational components that support our method: (1) the v-objective score-matching formulation for diffusion models [87], which underpins our latent modeling, and (2) the canonical inpainting algorithm introduced by Song et al. [11] and (3) further refined by Lugmayr et al. [89].
@@ A.3 RePaint
Lugmayr et al. proposed RePaint [89], a resampling-based mechanism that improves score-based inpainting by repeating the diffusion process across multiple forward–reverse cycles. Their key insight is that, in conventional inpainting (as described in Eq. (21)), the sampled noise for the known region is independent of the generated (inpainted) region. This lack of synchronization can lead to semantic inconsistencies and disharmony between the known and unknown parts of the sample.
@@ Appendix B Adaptive Timestep Conditioning and Score Approximation
where the approximation assumes that $\textbf{z}_{0}^{\Gamma}\sim q(\mathbf{x}_{0}^{\Gamma}|\textbf{z}_{0}^{\Omega})$ is available from the dataset. Unlike the marginalization-based approximation in Eq. (21), this expression introduces no sampling noise during inference, thereby reducing variance.
@@
$\nabla_{\textbf{z}_{\tau}^{\Gamma}}\log q_{\tau}(\textbf{z}_{\tau}^{\Gamma}|\textbf{z}_{0}^{\Omega})$ $\approx\frac{\alpha_{\tau}\textbf{z}_{0}^{\Gamma}-\textbf{z}_{\tau}^{\Gamma}}{\beta_{\tau}^{2}}$ (25)
$\approx\frac{\alpha_{\tau}\hat{\textbf{z}}_{\theta}(\textbf{z}_{\tau}^{\Gamma},\textbf{z}_{0}^{\Omega},\boldsymbol{\tau}^{\Gamma})-\textbf{z}_{\tau}^{\Gamma}}{\beta_{\tau}^{2}},$ (26)